A cyclist travels north along a road at a constant speed of 18 miles per hour. At 1:00 P.M., a runner is 44 miles away, running south along the same road at a constant speed. They pass each other at 3:00 P.M.. What is the speed of the runner?

Answer by ikleyn(52790)

.

    2*(18+x) = 44


    18 + x = 44/2 = 22


         x = 22 - 18 = 4 miles per hour.    ANSWER

Solved.

No explanations are necessary. Everything is as clear as the sun at the sunny day.